OpenStudy (anonymous):

The probability that a number is Divisible by 5 and more than 5000 can be made from digits 0,1,3,5,7 if repitation is not allowed(4 digit number)?

OpenStudy (anonymous):

4 Digits number more than 5000 can be formed from 0,1,3,5,7 = 2 * 4 * 3 *2 = 48 ways

OpenStudy (anonymous):

Divisible by 5 = 18 So p(Div 5) = 18 /48 .......am i correct

OpenStudy (anonymous):

 so 18/48 yup correct:)